OpenStudy (anonymous):

For the function f(x)= (8-2x)^2 , find f^-1. determine whether f^-1function

OpenStudy (goformit100):

solve it using limit

OpenStudy (anonymous):

Inverse...? \(\large\frac{1}{f(x)}\)

OpenStudy (anonymous):

Or do you mean the anti-function, as in anti-derivative for example?

OpenStudy (anonymous):

He is saying inverse function y=(8-2x)^2 x=(8-2y)^2 (+/-)(sqrt(x)-8)/2=y The inverse is not a function because there are 2 possible values for y

OpenStudy (anonymous):

FYI about being careful about notation: \[\tan^{-1} \theta = arctan \theta \neq \frac{1}{\tan \theta} = \cot \theta\] \[(x^2+1)^{-1} = \frac{1}{x^2+1}\]

OpenStudy (anonymous):

@ivan514 , If you assume that, then yes. I don't presume what he/she intended. :-)

OpenStudy (anonymous):

Statistically, most questions here are algebra2 level.. So most likely thats what he asked

OpenStudy (anonymous):

question is to find the inverse but it is not posed correctly because you cannot say "find the inverse function and then say if it is a function"

OpenStudy (anonymous):

this function is not one to one and so it does not have an inverse

OpenStudy (anonymous):

@ivan514 Meh... that's a bad assumption to make, it depends on the day and who's asking. As they say, "There's lies, damn lies, and statistics." lol I agree with @satellite73

OpenStudy (anonymous):

you can write \[x=(8-2y)^2\] \[\pm x=8-2y\] \[2y=8\pm\sqrt{x}\] \[y=4\pm\frac{\sqrt{x}}{2}\] if you like, but that does not make y a function of x

OpenStudy (anonymous):

typo on line two should be \[\pm\sqrt{x}=8-2y\]

OpenStudy (anonymous):

I understand its not a function but i its giving me 2 answer choices +/- = 8+/- sqrt x/2 and +/- sqrt 8+x /2

OpenStudy (anonymous):

Helpful syntax display tip: \pm \(\rightarrow\ \pm\) Put a \( before the \pm and

OpenStudy (anonymous):

a \) after it to activate

OpenStudy (anonymous):

\(8 \pm \sqrt{\large\frac{x}{2}}\) and \(\pm\sqrt{8 + \large\frac{x}{2}}\) , yes?

OpenStudy (anonymous):

Yes . These are my 2 answer choice.

OpenStudy (anonymous):

What I had: \[y=(8-2x)^2\] \[\pm \sqrt{y}=8-2x\] \[\pm \sqrt y - 8 = -2x\] \[\frac{\pm\sqrt{y}-8}{-2}=x\] \[\frac{8\pm\sqrt{y}}{2}=x\] \[4\pm \frac{\sqrt{y}}{2}=x\] switching x & y to make the switched around, anti-function : (the function that undoes what f(x) does) \[4\pm \frac{\sqrt{x}}{2}=y\] Same thing as @satellite73 had, he just switched the independent & dependent variables at the start, whereas I did it at the end. Are you sure it's not: f(x)=(16-2x)^2 ???

OpenStudy (anonymous):

This is also valid: \[\frac{8\pm \sqrt{x}}{2}=y\] But: \[\frac{8\pm \sqrt{x}}{2} \neq 8\pm \frac{\sqrt{x}}{2} \neq \pm\sqrt{8 + \frac{x}{2}}\]

OpenStudy (anonymous):

I final1y got the answer. Thanks so much you guys !